Explain why the results adhere to Ohm's Law


1
Expert's answer
2021-09-08T09:44:27-0400


We know that

Ohm's law

Potential difference "\alpha" electric current

"V\alpha i"

V=iR

Where

R =proportional constant



According to diagram

"\frac{1}{R}=\frac{∆ I}{∆V}\\R=\frac{∆V}{∆I}"